OpenStudy (***[isuru]***):

When A + B + C = pi prove that Sin 2A + sin 2B - sin 2C = 4cosAcosBcosC hence deduced that sin 2A + sin2B +sin 2C = 4sinAsinBsinC

OpenStudy (***[isuru]***):

Pls guys !!! Help me to solve this one!

OpenStudy (yttrium):

Use two product-to-sum formulas. \[sinAsinB = \frac{ 1 }{ 2 } [\cos(A-B) - \cos(A+B)]\] and \[cosAsinB = \frac{ 1 }{ 2 } [\sin(A+b) - \sin (A-B)]\]

OpenStudy (yttrium):

Let's try. \[4sinAsinBsinC = 4(\frac{ 1 }{ 2 } [\cos(A-B)-\cos(A+B)]) sinC\] \[=2\cos(A-B)sinC-2\cos(A+B)sinC\] \[2(\frac{ 1 }{ 2 }[\sin(A-B+C)-\sin(A-B-C)]) - 2(\frac{ 1 }{ 2}[\sin(A+B+C)-\sin(A+B-C)])\] \[=\sin(A-B+C)-\sin(A-B-C)-\sin(A+B+C)+\sin(A+B-C)\]